# Enhancing Conversational Agents from Open-Source Large Language Models with Illocutionary Force and Document-Based Knowledge Retrieval

### v1

@@ Abstract
In this paper, we first present a novel way of computationally analysing and extracting illocutionary forces from dialogue using Bert-based Large Language Models, and demonstrate how these features impact the response of a conversational agent guided by a document-based knowledge bank demonstrated by a bespoke web conversational chat agent system developed. Our proposed illocutionary force extraction and classification technique is the first of its kind using the Argument Interchange Format (AIF) Dataset, showing an improved performance compared to two methods for carrying out similar tasks with a macro F1 of approximately 45%. When we evaluated the system based on 2 knowledge files, with 2 user queries each, across 5 open-source large language models (LLMs) using 10 standard metrics we found out that larger open-source models, such as Llama2:13b and Llama3-chatqa-latest, demonstrated an improved alignment when the user illocutionary force was included with their query, achieving higher QA and linguistic similarity scores. The smaller models on the other hand like Tinyllama:latest showed an increased perplexity and mixed performance, which explicitly indicated struggles in processing queries that explicitly included illocutionary forces. The results from the analysis highlight the potential of illocutionary force to enhance conversational depth while underscoring the need for model-specific optimizations to address increased computational costs and response times.
@@ 1 Introduction
The spontaneous evolution and transformation of conversational agents, ignited by the improvements in LLMs has transformed the way in which humans interact and respond with technology. Due to factors like adaptability, accessibility, and performance, open-source LLMs have gained significant attention. Regardless of these strengths, these open-source models often struggle with refined linguistic features such as illocutionary force which forms the the intended speech act behind a statement or locution, creating a gap in better understanding the speaker or user’s communicative intent. Handling this limitation is essential for improving the relevance, coherence, and responsiveness of conversational agents.
This research work introduces a novel framework for analyzing and extracting illocutionary forces behind speakers’ or authors’ locutions using BERT-based transformer architectures and integrating these extracted features into a custom document-based conversational system. When we combined the AIF dataset with our novel dynamic class weight balance technique to tackle class imbalance, we achieved a better macro F1 score which outperformed existing methods by approximately 10% to 15%. Additionally, this research work explores the impact of integrating the user illocutionary force into the user queries for the conversational agents which leverages document-based knowledge retrieval to improve the quality of the response and effectively evaluate its performance using standard metrics which cross-reference the chat agent response to a pre-existing answer stored somewhere (information from a custom file). The experiments we conducted showed varying performance across the 5 open-source models used across the 11 standard evaluation metrics used. The impacts of this research are twofold:
@@
    It exposes the normally overlooked role and impact of illocutionary force in discourse analysis and its broad impact on conversational agents.
@@ 2 Literature Review
It has been seen that open-source LLMs such as LLaMA [17], Mistral [18], and TinyLLama among others have risen as a relevant and needful alternative to closed-source models because of the scalability, flexibility, and creativity in which they offer. These open-source models have demonstrated imminent relevance in various NLP and machine learning tasks such as natural language generation (NLG) and natural language understanding (NLU). While praising these models for their remarkable ability to capture semantic and syntactic nuances in various tasks, they still struggle with capturing pragmatic nuances such as illocutionary force [6]. Addressing this limitation opens up a less trivial opportunity for additional linguistic and contextual analysis interfaces to improve conversational applications.
@@
To give a background Illocutionary Force in Text Analysis, Illocutionary force is often referred to as the communicative intent behind a statement or locution, as defined in speech act theory [1] [15]. Thus far, there has been a lot of visible significance of the concept of illocutionary forces theoretically, but very limited practical application to NLP. It is seen that more focused text analysis for sentiment, emotion, and intent classification has been enabled by recent improvements in transformer-based architecture models, such as BERT and its derivatives [5]. However, illocutionary force classification remains underexplored and overlooked in computational studies. The core of this study is to build on existing foundations of the speech act theory, leveraging the AIF dataset to model illocutionary forces in dialogue structures, thereby filling in a critical research gap in text analysis for conversational systems.
3 Illocutionary Force Modelling
3.1 Dataset
To give a summary of the data structure and components of the data set worked with, we made use of various annotated corpus stored in the Argument Interchange Format Database (AIFdb) [8] with a JSON format. The key data components in the structure are the nodes which represent various discourse-argumentative units, edges which show how various nodes are inter-connected, and the locution nodes which give details about the participants in the discourse. The Corpora used were US2016, QT30, MM123, and MM2012. We represent these four corpora with ‘UQMM’. When analyzing the distribution of the illocutionary forces we first had all these class lists:
Original_Labels = [
    ’Asserting’, ’Disagreeing’, ’Pure Questioning’, ’Agreeing’,
    ’Assertive Questioning’, ’Rhetorical Questioning’, ’Popular’,
    ’Default Illocuting’, ’Arguing’, ’Assertive Challenging’,
    ’Challenging’, ’Ironic Asserting’, ’Rhetorical Challenging’,
    ’Offering’, ’Pure Challenging’, ’Restating’, ’Ad verecundiam’,
    ’Directive Questioning’, ’Order’, ’Conceding’, ’YA’, ’Expressing’,
    ’Analysing’, ’Conceding’
Following a Qualitative analysis, we pruned these labels [“YA”, “Expressing”, “Offering”, “Ad verecundiam”, “Order”, “Analysing”, “Conceding”, “Popular Conceding”], as some were errors from annotation and others were not needed.
At the final analysis stage, we noticed that some class labels were closely related, we then decided to merge closely related classes together to minimize bias during machine learning. These were “Directive Questioning into Pure Questioning”, “Ironic Asserting into Asserting”, and “Assertive Challenging, Pure Challenging, Rhetorical Challenging into Challenging”. We show three lists below which consist of the class lists after the running and merging pre-processing task.
Original_Labels = [
    ’Asserting’, ’Disagreeing’, ’Pure Questioning’, ’Agreeing’,
    ’Assertive Questioning’, ’Rhetorical Questioning’, ’Popular’,
    ’Default Illocuting’, ’Arguing’, ’Assertive Challenging’,
    ’Challenging’, ’Ironic Asserting’, ’Rhetorical Challenging’,
    ’Offering’, ’Pure Challenging’, ’Restating’, ’Ad verecundiam’,
    ’Directive Questioning’, ’Order’, ’Conceding’, ’YA’, ’Expressing’,
    ’Analysing’, ’Conceding’
Prunning_Results = [
    ’Asserting’, ’Disagreeing’, ’Pure Questioning’, ’Agreeing’,
    ’Assertive Questioning’, ’Rhetorical Questioning’, ’Conceding’,
    ’Default Illocuting’, ’Arguing’, ’Assertive Challenging’,
    ’Challenging’, ’Ironic Asserting’, ’Rhetorical Challenging’,
    ’Challenging’, ’Pure Challenging’, ’Restating’,
    ’Directive Questioning’, ’Pure Questioning’
Merged_Labels = [
    ’Asserting’, ’Disagreeing’, ’Pure Questioning’, ’Agreeing’,
    ’Assertive Questioning’, ’Rhetorical Questioning’,
    ’Default Illocuting’, ’Arguing’, ’Challenging’,
    ’Challenging’, ’Asserting’, ’Challenging’,
    ’Challenging’, ’Restating’
Table 2 below further shows the class distributions across the corpus combinations experimented with.
Prunned Labels UQ UQMM Merged Labels UQMM
Agreeing 108 110 Agreeing 110
Arguing 33 33 Arguing 33
Asserting 13820 27449 Asserting 27476
Assertive Challenging 38 45 Assertive Questioning 629
Assertive Questioning 493 629 Challenging 118
Challenging 42 42 Default Illocuting 68
Conceding 2 4 Disagreeing 26
Default Illocuting 66 68 Pure Questioning 1471
Disagreeing 26 26 Restating 10
Ironic Asserting 26 27 Rhetorical Questioning 404
Popular Conceding 2 30
Pure Challenging 10 15
Pure Questioning 1382 1471
Restating 5 10
Rhetorical Challenging 3 16
Rhetorical Questioning 313 404
Table 1: Class Counts
3.2 Pre-processing
Our preprocessing stage was divided into four major stages seen below:
  • Corpus files concatenation: Each file constituent of the corpus stored in the AIFdb was merged into one large file consisting of all the node elements identifiers (nodeID, text, type, timestamp) and unique values in one list with the key denoted as “nodes”, and all the edges elements identifiers (edgeID, fromID, toID) and values in another list with the key denoted as “edges”.
  • Features extraction: At this stage, we attempt to extract all the features (Locutions and Illocutionary forces) by looping through the data to identify all the locutions nodes denoted by ‘L’ which is a fromID and the toID is an illocutionary force node denoted by ‘YA’. We defined the fromID key in the edges list as ‘from_id’, and the toID key as ‘to_id’. Finally, we iterate through all the edges, and for each fromID which had a type value of L in the nodes list, and the toID of the same edge map with a type value of YA in the nodes list, we add the L text value, and the YA text value to an initialized list of Locutions and Illocutionary Forces while ensuring that their indexes match. This iteration continues until all the conditions are met through the data.
  • Speaker name stripping: After deriving all the locutions and illocutionary forces linked together, at the stage, we progressed to using a regex match to strip out the Speaker names from the locutions which could introduce noise during machine learning.
  • Data restructuring: We finally generated a resultant file which was the training data to be used. These were the locutions and illocutionary forces from the initialized lists with their from and to ID values appended into various JSON maps with a list denoted by illocutions.
    Figure 1: Training Data Structure
3.3 Model Fine-tuning
Using an inverse frequency fine-tuning novel approach of transformer models for handling class imbalance [11], we fine-tuned a pre-trained Bert-base, Distilbert, Roberta-base, and Roberta-large model on an 80:20 train-test split with learning_rate=2e-5, epoch=20, batch_size=50, max_length=50, seed=42, and using cross-entropy loss function. The ‘L_text’ key value consisting of the author utterance from the result of the preprocessing served as the model input while the ‘YA_text’ key value consisting of the illocutionary forces served as the class labels.
Corpus Model Precision Recall F1 Score Accuracy
UQMM Fine-tuned Bert-base 0.43 0.37 0.39 0.64
Fine-tuned Distilbert 0.39 0.34 0.36 0.63
Fine-tuned RoBERTa-base 0.59 0.48 0.52 0.65
Fine-tuned RoBERTa-large 0.57 0.52 0.54 0.66
Table 2: Performance across Transformer Models on Standard Approaches
3.4 Comparison and Evaluation
Comparing our fine-tuning modelling approach with other methods, we further explored two other methods which have been used to address an identical task to ours. These are:
  • Extreme multi-label text classification using CNN: This approach was used for classifying multiple labels from very large label classes to various documents using Convolutional Neural Networks (CNN) forming the core of their work on Extreme multi-label text classification (XMTC) using CNN [10].
  • Few-Shot Using Transformers with Low-Rank Adaptation: Very recently, Jairam et al. [12] proposed an approach which incorporated transformer models fine-tuned and refined with Low-Rank Adaptation of Large Language Models (LoRA) and Quantized Low-Rank Adaptation of Large Language Models (QLoRA) with few-shot learning on six Indian English accents in the Indic Accent imbalanced Dataset with distribution of (Tamil at 1,640, Malayalam at 1,563, 175 Telugu at 1,614, Gujarati at 298, Hindi at 827, Kannada at 1,486) which showed an Accuracy of up to 94% with trainable parameters set at 5%. We only compared with LoRA for this research.
The table below shows the performance of our approach as compared to the two approaches above.
Method Model Epoch Precision Recall F1 Score Accuracy
Inverse Class Frequency Fine-tuned Bert-base 20 0.43 0.37 0.39 0.64
Fine-tuned Distilbert 20 0.39 0.34 0.36 0.63
Fine-tuned RoBERTa-base 20 0.59 0.48 0.52 0.65
Fine-tuned RoBERTa-large 20 0.57 0.52 0.54 0.66
Other Methods XMTC CNN 10 0.25 0.26 0.25 0.50
BERT + LoRA 10 0.04 0.01 0.06 0.38
BERT + LoRA 20 0.08 0.05 0.16 0.44
BERT + LoRA 40 0.18 0.21 0.19 0.63
RoBERTa + LoRA 10 0.12 0.14 0.12 0.47
RoBERTa + LoRA 20 0.20 0.21 0.19 0.57
RoBERTa + LoRA 40 0.35 0.36 0.35 0.69
Table 3: Performance across methods
@@ 4 System Overview
The Web Application in Figure 2 was custom-developed for the purpose of this research. It offers the public the ability to carry out research and experiments with illocutionary forces, knowledge base, and open-source LLMs. We give an overview of three key features used for this research below:
@@ 4.1 Model Selection
4.2 Illocutionary Force Identification
This component forms the backbone of this research as it gives the ability to either include the user illocutionary force from their query alongside their query to the conversational agent or just pass their query alone to the conversational agent to generate a response.
@@ 4.3 Document-Based Knowledge Retrieval
In this integration, we aimed to incorporate illocutionary force into response generation from information during document retrieval for enhancing contextual alignment between the retrieved information and the user intent. The end users have the ability to upload and select various knowledge files in either .txt or .pdf formats to use as the conversational agent’s knowledge to guide its response.
@@ 5 Experiment Design
To fully understand and expose the impacts of including the illocutionary force of a user query alongside their query, we set up two key experiments.
Experiment 1: In the first experiment, we used two custom documents seen in section 5.1 below as a knowledge base of the model to restrict its knowledge to only the information in the document. We provided two queries for each of the 5 models while assessing without including the users’ illocutionary force, and evaluated the performance of each model based on 10 metrics seen in section 5.5.
@@
Experiment 2: In the second experiment, we used two custom documents seen in section 5.1 below as a knowledge base of the model to restrict its knowledge to only the information in the document. We provided two queries and the illocutionary forces extracted from the queries for each of the 5 models and evaluated the performance of each model based on 10 metrics seen in section 5.5.
User Query + User Illocutionary ForceConversational AgentOllama ModelCustom DocumentGenerated Response
The User Query and Illocutionary Force serves as the input to the conversational AI system. This conversational AI system is driven by 5 Ollama Model which is guided by Custom Documents to produce the required Generated Response based on the prompt designed.
@@ 5.2 Prompt Design
In designing the prompt for the conversational agent, four parameters were used. The ‘user_input’ takes the user query argument from the front-end; The ‘context’ consists of a list of the previous conversation between the user and the conversational agent with indicators of users and assistants so that the agent can understand which previous conversation was the user and which was from itself just to ensure coherence; The ‘knowledge_text’ which consist of the text content of the knowledge document selected from the front-end, and the ‘illocutionary_force’ which consist of the illocutionary force class assigned to the user query after inference from the model stored on hugging-face. The inclusion of the ‘illocutiomnary_force’ is been controlled from the front-end hence the condition in the prompt which is initialised to ‘True’ but updates based on the choice of the end-user. We show these representations below in a pseudo-code.
@@
instruction += f"Consider the user’s communicative intent while responding, characterized by the illocutionary force: ’{illocutionary_force}’."
@@ 5.4 Conversational agent response evaluation metrics
  • Recall-Oriented Understudy for Gisting Evaluation (ROUGE-1): Rouge is a conversation metric which is used to evaluate the quality of a response (in most cases - a chatbot) comparing its overlapping unigram with a standardise reference response (normally a human-generated response) [9]. In this work, ROUGE-1 is used to measure the presence of certain keywords in the generated response of the chatbot with reference to the custom document used to guide its response. A high ROUGE-1 score indicates that the chatbot’s response captures the necessary relevant information from the reference document.
@@
  • ROUGE-L: ROUGE-L on the other hand evaluates the longest common subsequence between the generated chatbot response and the custom document been used to guide its response. It covers a sentence level, unlike ROUGE-1 and ROUGE-2, making it a good evaluator for syntactic and discourse-level performance. Lin et al., in 2004 [9] argued that ROUGE-L was better for capturing sentence level coherence as compared with ROUGE-1 and ROUGE-2. A higher ROUGE-L score indicates that the chatbot response maintains a similar structure as compared to the reference custom document.
@@ 6 Experimental Results and Analysis
We first installed OLLAMA SDK and pulled the 5 models we needed for our work to our local computer. Our computer had a GPU of 12 GB RAM NVIDIA GEFORCE. After following the experimental set-up in section 5 above, we recorded the results from all the metrics used from the custom-designed web application including the response time after the query was passed to the model. We show the results in the two tables below with the first one consisting of the results without including the illocutionary force of the query, and the second table consisting of the results after the illocutionary force of the query was included.
Without Illocutionary Force added to User Query:
S/N Documents Model Bot Response BERT BERT BERT QA QA ROUGE-1 ROUGE-2 ROUGE-L METEOR Perplexity Response Time (sec) Precision Recall F1 Ref Cand 1 001 llama2:13b First 74.64 0.72 0.72 0.72 0.49 0.24 0.24 0.13 0.19 0.42 56.75 2 001 Tinyllama:latest First 26.55 0.68 0.68 0.68 0.49 0.39 0.18 0.07 0.12 0.35 26.69 3 001 Llama3-chatqa-latest First 74.33 0.84 0.84 0.84 0.49 0.37 0.43 0.21 0.38 0.49 54.7 4 001 Llama3-latest First 105.02 0.74 0.74 0.74 0.49 0.63 0.16 0.01 0.09 0.22 49.16 5 001 Mistral:latest First 109.38 0.57 0.57 0.57 0.49 0.55 0.19 0.04 0.12 0.30 28.64 6 001 llama2:13b Second 164.58 0.92 0.92 0.92 0.04 0.17 0.39 0.13 0.24 0.46 70.62 7 001 Tinyllama:latest Second 18.04 0.81 0.81 0.81 0.04 0.09 0.33 0.09 0.18 0.41 48.82 8 001 Llama3-chatqa-latest Second 62.80 0.90 0.90 0.90 0.04 0.31 0.25 0.08 0.18 0.26 29.12 9 001 Llama3-latest Second 85.69 0.86 0.86 0.86 0.04 0.17 0.19 0.04 0.13 0.28 55.14 10 001 Mistral:latest Second 130.64 0.79 0.79 0.79 0.04 -0.03 0.17 0.06 0.10 0.31 28.96 11 004 llama2:13b First 543.83 0.78 0.78 0.78 0.77 0.63 0.12 0.05 0.10 0.21 41.60 12 004 Tinyllama:latest First 57.02 0.84 0.84 0.84 0.77 0.03 0.10 0.00 0.08 0.15 1849.46 13 004 Llama3-chatqa-latest First - - - - - - - - - - - 14 004 Llama3-latest First 252.28 0.74 0.74 0.74 0.77 0.68 0.11 0.04 0.08 0.19 26.80 15 004 Mistral:latest First 279.70 0.80 0.80 0.80 0.77 0.65 0.14 0.06 0.11 0.22 29.15 16 004 llama2:13b Second 532.95 0.89 0.89 0.89 0.02 -0.01 0.15 0.09 0.11 0.40 45.45 17 004 Tinyllama:latest Second 52.84 0.93 0.93 0.93 0.02 0.03 0.19 0.05 0.19 0.20 2612.20 18 004 Llama3-chatqa-latest Second - - - - - - - - - - - 19 004 Llama3-latest Second 237.64 0.89 0.89 0.89 0.02 0.10 0.25 0.05 0.16 0.29 35.32 20 004 Mistral:latest Second 298.41 0.82 0.82 0.82 0.02 0.07 0.15 0.06 0.10 0.23 34.43
With Illocutionary Force added to User Query:
S/N Documents Model Bot Response BERT BERT BERT QA QA ROUGE-1 ROUGE-2 ROUGE-L METEOR Perplexity Response Time (sec) Precision Recall F1 Ref Cand 1 001 llama2:13b First 212.15 0.79 0.79 0.79 0.49 0.51 0.22 0.07 0.12 0.33 72.52 2 001 Tinyllama:latest First 24 0.74 0.74 0.74 0.49 0.48 0.23 0.10 0.13 0.41 32.90 3 001 Llama3-chatqa-latest First 70.92 0.64 0.64 0.64 0.49 0.68 0.24 0.02 0.14 0.24 58.74 4 001 Llama3-latest First 92.45 0.66 0.66 0.66 0.49 0.63 0.19 0.02 0.11 0.28 60.24 5 001 Mistral:latest First 107.49 0.73 0.73 0.73 0.49 0.45 0.24 0.07 0.13 0.37 28.67 6 001 llama2:13b Second 244.50 0.79 0.79 0.79 0.043 0.066 0.22 0.07 0.12 0.33 73.07 7 001 Tinyllama:latest Second 28.60 0.69 0.69 0.69 0.04 0.03 0.17 0.07 0.10 0.29 29.29 8 001 Llama3-chatqa-latest Second 64.05 0.98 0.98 0.98 0.07 0.60 0.14 0.0 0.14 0.10 70.82 9 001 Llama3-latest Second 80.86 0.87 0.87 0.87 0.04 0.11 0.23 0.03 0.15 0.32 20.44 10 001 Mistral:latest Second 108.78 0.83 0.83 0.83 0.04 0.09 0.25 0.06 0.17 0.37 32.75 11 004 llama2:13b First 611.63 0.83 0.83 0.83 0.77 0.69 0.13 0.05 0.09 0.25 39.09 12 004 Tinyllama:latest First 59.09 0.79 0.79 0.79 0.77 0.15 0.13 0.03 0.09 0.16 2552.25 13 004 Llama3-chatqa-latest First - - - - - - - - - - - 14 004 Llama3-latest First 350.84 0.80 0.80 0.80 0.77 0.64 0.11 0.03 0.07 0.19 31.40 15 004 Mistral:latest First 375.14 0.76 0.76 0.76 0.77 0.61 0.12 0.04 0.09 0.28 51.88 16 004 llama2:13b Second 613.46 0.82 0.82 0.82 0.02 -0.067 0.15 0.05 0.08 0.22 51.75 17 004 Tinyllama:latest Second 67.30 0.86 0.86 0.86 0.77 0.03 0.09 0.05 0.06 0.25 583.08 18 004 Llama3-chatqa-latest Second - - - - - - - - - - - 19 004 Llama3-latest Second 330.94 0.86 0.86 0.86 0.019 0.017 0.21 0.042 0.12 0.28 104.75 20 004 Mistral:latest Second 329.08 0.83 0.83 0.83 0.02 -0.019 0.20 0.07 0.12 0.32 41.03

### v2

# An Open-Source Web-Based Tool for Evaluating Open-Source Large Language Models Leveraging Information Retrieval from Custom Documents

@@ Abstract
In our work, we present the first-of-its-kind open-source web-based tool which is able to demonstrate the impacts of a user’s speech act during discourse with conversational agents, which leverages open-source large language models. With this software resource, it is possible for researchers and experts to evaluate the performance of various dialogues, visualize the user’s communicative intents, and utilise uploaded specific documents for the chat agent to use for its information retrieval to respond to the user query. Almost all modern conversational agents and systems leverage the encoder-decoder architecture of generative models to first understand and generate responses to user queries. The context gathered by these models is obtained from a set of linguistic features extracted, which forms the context embeddings of the models. Regardless of these models showing good context understanding based on these features, there still remains a gap in including deeper pragmatic features to improve the model’s comprehension of the query, hence the efforts to develop this web resource, which is able to extract and then inject this overlooked feature in the encoder-decoder pipeline of the conversational agent. To demonstrate the effect and impact of the resource, we carried out an experiment which evaluated the system using 2 knowledge files for information retrieval, with 2 user queries each, across 5 open-source large language models using 10 standard metrics. Our results showed that larger open-source models, such as (e.g., Llama2:13B, Llama3-ChatQA-Latest) demonstrated an improved alignment when the user speech act was included with their query, achieving higher Question-Answer and linguistic similarity scores to the source document. The smaller models in contrast (e.g., TinyLlama:Latest) showed an increased perplexity and mixed performance, which explicitly indicated struggles in processing queries that explicitly included speech act. The results from the analysis using the developed web resource highlight the potential of speech acts towards enhancing conversational depths while underscoring the need for model-specific optimizations to address increased computational costs and response times.
@@ 1 Introduction
The spontaneous evolution and transformation of conversational agents, ignited by the improvements in LLMs has transformed the way in which humans interact and respond with technology. Due to factors like adaptability, accessibility, and performance, open-source LLMs have gained significant attention. Regardless of these strengths, these open-source models often struggle with refined linguistic features such as speech acts which forms the the intended speech act behind a statement or locution, creating a gap in better understanding the speaker or user’s communicative intent. Handling this limitation is essential for improving the relevance, coherence, and responsiveness of conversational agents.
Additionally, this research work explores the impact of integrating the user speech acts into the user queries for the conversational agents which leverages document-based knowledge retrieval to improve the quality of the response and effectively evaluate its performance using standard metrics which cross-reference the chat agent response to a pre-existing answer stored somewhere (information from a custom file). The experiments we conducted showed varying performance across the 5 open-source models used across the 11 standard evaluation metrics used. The impacts of this research are twofold:
@@
    It exposes the impact of speech acts in discourse analysis and its broad impact on conversational agents.
@@ 2 Literature Review
It has been seen that open-source LLMs such as LLaMA [12], Mistral [13], and TinyLLama among others have risen as a relevant and needful alternative to closed-source models because of the scalability, flexibility, and creativity in which they offer. These open-source models have demonstrated imminent relevance in various NLP and machine learning tasks such as natural language generation (NLG) and natural language understanding (NLU). While praising these models for their remarkable ability to capture semantic and syntactic nuances in various tasks, they still struggle with capturing pragmatic nuances such as speech acts [6]. Addressing this limitation opens up a less trivial opportunity for additional linguistic and contextual analysis interfaces to improve conversational applications.
@@
To give a background of speech acts in Text Analysis, speech acts are often referred to as the communicative intent behind a statement or locution, as defined in speech act theory [1] [10]. Thus far, there has been a lot of visible significance of the concept of speech acts theoretically, but very limited practical application to NLP. It is seen that more focused text analysis for sentiment, emotion, and intent classification has been enabled by recent improvements in transformer-based architecture models, such as BERT and its derivatives [5]. However, speech act classification remains underexplored and overlooked in computational studies. The core of this study is to build on existing foundations of the speech act theory, leveraging the AIF dataset to model speech acts in dialogue structures, thereby filling in a critical research gap in text analysis for conversational systems.
@@ 3 System Overview
The Web Application in Figure 1 was custom-developed for the purpose of this research. It offers the public the ability to carry out research and experiments with speech actss, knowledge base, and open-source LLMs. We give an overview of three key features used for this research below:
@@ 3.1 Model Selection
3.2 Speech Acts Identification
This component forms the backbone of this research as it gives the ability to either include the user speech acts from their query alongside their query to the conversational agent or just pass their query alone to the conversational agent to generate a response. The speech act identification uses the model stored here  [^1] Model.
@@ 3.3 Document-Based Knowledge Retrieval
In this integration, we aimed to incorporate speech acts into response generation from information during document retrieval for enhancing contextual alignment between the retrieved information and the user intent. The end users have the ability to upload and select various knowledge files in either .txt or .pdf formats to use as the conversational agent’s knowledge to guide its response.
@@ 4 Experiment Design
To fully understand and expose the impacts of including the speech acts of a user query alongside their query, we set up two key experiments.
Experiment 1: In the first experiment, we used two custom documents seen in section 5.1 below as a knowledge base of the model to restrict its knowledge to only the information in the document. We provided two queries for each of the 5 models while assessing without including the users’ speech acts, and evaluated the performance of each model based on 10 metrics seen in section 5.5.
@@
Experiment 2: In the second experiment, we used two custom documents seen in section 5.1 below as a knowledge base of the model to restrict its knowledge to only the information in the document. We provided two queries and the speech actss extracted from the queries for each of the 5 models and evaluated the performance of each model based on 10 metrics seen in section 5.5.
User Query + User speech actsConversational AgentOllama ModelCustom DocumentGenerated Response
The User Query and speech acts serves as the input to the conversational AI system. This conversational AI system is driven by 5 Ollama Model which is guided by Custom Documents to produce the required Generated Response based on the prompt designed.
@@ 4.2 Prompt Design
In designing the prompt for the conversational agent, four parameters were used. The ‘user_input’ takes the user query argument from the front-end; The ‘context’ consists of a list of the previous conversation between the user and the conversational agent with indicators of users and assistants so that the agent can understand which previous conversation was the user and which was from itself just to ensure coherence; The ‘knowledge_text’ which consist of the text content of the knowledge document selected from the front-end, and the ‘illocutionary_force’ which consist of the speech acts class assigned to the user query after inference from the model stored on hugging-face. The inclusion of the ‘illocutiomnary_force’ is been controlled from the front-end hence the condition in the prompt which is initialised to ‘True’ but updates based on the choice of the end-user. We show these representations below in a pseudo-code.
@@
instruction += f"Consider the user’s communicative intent while responding, characterized by the speech acts: ’{illocutionary_force}’."
@@ 4.4 Conversational agent response evaluation metrics
  • Recall-Oriented Understudy for Gisting Evaluation (ROUGE-1): Rouge is a conversation metric which is used to evaluate the quality of a response (in most cases - a chatbot) comparing its overlapping unigram with a standardise reference response (normally a human-generated response) [lin-2002-rouge]. In this work, ROUGE-1 is used to measure the presence of certain keywords in the generated response of the chatbot with reference to the custom document used to guide its response. A high ROUGE-1 score indicates that the chatbot’s response captures the necessary relevant information from the reference document.
@@
  • ROUGE-L: ROUGE-L on the other hand evaluates the longest common subsequence between the generated chatbot response and the custom document been used to guide its response. It covers a sentence level, unlike ROUGE-1 and ROUGE-2, making it a good evaluator for syntactic and discourse-level performance. Lin et al., in 2002 [lin-2002-rouge] argued that ROUGE-L was better for capturing sentence level coherence as compared with ROUGE-1 and ROUGE-2. A higher ROUGE-L score indicates that the chatbot response maintains a similar structure as compared to the reference custom document.
@@ 5 Experimental Results and Analysis
We first installed OLLAMA SDK and pulled the 5 models we needed for our work to our local computer. Our computer had a GPU of 12 GB RAM NVIDIA GEFORCE. After following the experimental set-up in section 5 above, we recorded the results from all the metrics used from the custom-designed web application including the response time after the query was passed to the model. We show the results in the two tables below with the first one consisting of the results without including the speech acts of the query, and the second table consisting of the results after the speech acts of the query was included.
Without speech acts added to User Query:
Table 1: Without speech acts added to User Query:
S/N Documents Model Bot Response BERT BERT BERT QA QA ROUGE-1 ROUGE-2 ROUGE-L METEOR Perplexity
Response Time (sec) Precision Recall F1 Ref Cand
1 001 llama2:13b First 74.64 0.72 0.72 0.72 0.49 0.24 0.24 0.13 0.19 0.42 56.75
2 001 Tinyllama:latest First 26.55 0.68 0.68 0.68 0.49 0.39 0.18 0.07 0.12 0.35 26.69
3 001 Llama3-chatqa-latest First 74.33 0.84 0.84 0.84 0.49 0.37 0.43 0.21 0.38 0.49 54.7
4 001 Llama3-latest First 105.02 0.74 0.74 0.74 0.49 0.63 0.16 0.01 0.09 0.22 49.16
5 001 Mistral:latest First 109.38 0.57 0.57 0.57 0.49 0.55 0.19 0.04 0.12 0.30 28.64
6 001 llama2:13b Second 164.58 0.92 0.92 0.92 0.04 0.17 0.39 0.13 0.24 0.46 70.62
7 001 Tinyllama:latest Second 18.04 0.81 0.81 0.81 0.04 0.09 0.33 0.09 0.18 0.41 48.82
8 001 Llama3-chatqa-latest Second 62.80 0.90 0.90 0.90 0.04 0.31 0.25 0.08 0.18 0.26 29.12
9 001 Llama3-latest Second 85.69 0.86 0.86 0.86 0.04 0.17 0.19 0.04 0.13 0.28 55.14
10 001 Mistral:latest Second 130.64 0.79 0.79 0.79 0.04 -0.03 0.17 0.06 0.10 0.31 28.96
11 002 llama2:13b First 543.83 0.78 0.78 0.78 0.77 0.63 0.12 0.05 0.10 0.21 41.60
12 002 Tinyllama:latest First 57.02 0.84 0.84 0.84 0.77 0.03 0.10 0.00 0.08 0.15 1849.46
13 002 Llama3-chatqa-latest First - - - - - - - - - - -
14 002 Llama3-latest First 252.28 0.74 0.74 0.74 0.77 0.68 0.11 0.04 0.08 0.19 26.80
15 002 Mistral:latest First 279.70 0.80 0.80 0.80 0.77 0.65 0.14 0.06 0.11 0.22 29.15
16 002 llama2:13b Second 532.95 0.89 0.89 0.89 0.02 -0.01 0.15 0.09 0.11 0.40 45.45
17 002 Tinyllama:latest Second 52.84 0.93 0.93 0.93 0.02 0.03 0.19 0.05 0.19 0.20 2612.20
18 002 Llama3-chatqa-latest Second - - - - - - - - - - -
19 002 Llama3-latest Second 237.64 0.89 0.89 0.89 0.02 0.10 0.25 0.05 0.16 0.29 35.32
20 002 Mistral:latest Second 298.41 0.82 0.82 0.82 0.02 0.07 0.15 0.06 0.10 0.23 34.43
With speech acts added to User Query:
Table 2: With speech acts added to User Query:
S/N Documents Model Bot Response BERT BERT BERT QA QA ROUGE-1 ROUGE-2 ROUGE-L METEOR Perplexity
Response Time (sec) Precision Recall F1 Ref Cand
1 001 llama2:13b First 212.15 0.79 0.79 0.79 0.49 0.51 0.22 0.07 0.12 0.33 72.52
2 001 Tinyllama:latest First 24 0.74 0.74 0.74 0.49 0.48 0.23 0.10 0.13 0.41 32.90
3 001 Llama3-chatqa-latest First 70.92 0.64 0.64 0.64 0.49 0.68 0.24 0.02 0.14 0.24 58.74
4 001 Llama3-latest First 92.45 0.66 0.66 0.66 0.49 0.63 0.19 0.02 0.11 0.28 60.24
5 001 Mistral:latest First 107.49 0.73 0.73 0.73 0.49 0.45 0.24 0.07 0.13 0.37 28.67
6 001 llama2:13b Second 244.50 0.79 0.79 0.79 0.043 0.066 0.22 0.07 0.12 0.33 73.07
7 001 Tinyllama:latest Second 28.60 0.69 0.69 0.69 0.04 0.03 0.17 0.07 0.10 0.29 29.29
8 001 Llama3-chatqa-latest Second 64.05 0.98 0.98 0.98 0.07 0.60 0.14 0.0 0.14 0.10 70.82
9 001 Llama3-latest Second 80.86 0.87 0.87 0.87 0.04 0.11 0.23 0.03 0.15 0.32 20.44
10 001 Mistral:latest Second 108.78 0.83 0.83 0.83 0.04 0.09 0.25 0.06 0.17 0.37 32.75
11 002 llama2:13b First 611.63 0.83 0.83 0.83 0.77 0.69 0.13 0.05 0.09 0.25 39.09
12 002 Tinyllama:latest First 59.09 0.79 0.79 0.79 0.77 0.15 0.13 0.03 0.09 0.16 2552.25
13 002 Llama3-chatqa-latest First - - - - - - - - - - -
14 002 Llama3-latest First 350.84 0.80 0.80 0.80 0.77 0.64 0.11 0.03 0.07 0.19 31.40
15 002 Mistral:latest First 375.14 0.76 0.76 0.76 0.77 0.61 0.12 0.04 0.09 0.28 51.88
16 002 llama2:13b Second 613.46 0.82 0.82 0.82 0.02 -0.067 0.15 0.05 0.08 0.22 51.75
17 002 Tinyllama:latest Second 67.30 0.86 0.86 0.86 0.77 0.03 0.09 0.05 0.06 0.25 583.08
18 002 Llama3-chatqa-latest Second - - - - - - - - - - -
19 002 Llama3-latest Second 330.94 0.86 0.86 0.86 0.019 0.017 0.21 0.042 0.12 0.28 104.75
20 002 Mistral:latest Second 329.08 0.83 0.83 0.83 0.02 -0.019 0.20 0.07 0.12 0.32 41.03